OpenStudy (anonymous):

6z-8-z+3=0? frist step is to combine the like tearms? so i would combime the z, but what about the negitive sign?

OpenStudy (anonymous):

We define subtraction as: \[ a-b=a+(-b) \]So, try using that.

OpenStudy (anonymous):

so the z would be like -1z?

OpenStudy (anonymous):

Yep, so now we can move it around in the equation, to see: \[ 6z-8-z+3=6z-8+(-z)+3=\\6z+(-z)-8+3=6z-z-8+3=5z-8+3=0 \]Et al.